# VFMM3D: Releasing the Potential of Image by Vision Foundation Model for Monocular 3D Object Detection

### v1

@@ IV-B Comparison with SOTA methods
We compare the performance of our method on KITTI val set for the car class with the state-of-the-art methods. The related results are shown in Tab. I. It is evident that our method outperforms existing methods in both 3D and BEV under all difficulty levels. Specifically, when IoU is set to 0.5, our method surpasses these methods by even greater margins. Compared to the latest monocular object detection method NeurOCS [47], our approach surpasses it by 3.36% and 5.02% in 3D and BEV at the easy level, respectively. Some qualitative results of 3D detection boxes from our method on KITTI val set are visualized in Fig. 4.
IV-C Visualization of Each Component Results.
@@ IV-D Ablation Study
Impacts of Different Painting Strategy. We conduct ablation experiments to assess the impacts of different pseudo-LiDAR painting strategies on the final 3D detection results. As depicted in Tab. II, we observe that the detailed and accurate foreground information provided by the text-prompt SAM (SAM-P) enables subsequent LiDAR-based 3D detectors to efficiently concentrate on the features of foreground objects. In comparison to the global painting approach (RGB-P), this enhances the overall average detection accuracy by 3% and 3.13% for 3D and BEV detection results, respectively.
Ablation Study on Different LiDAR-Based 3D Object Detectors. We employ various LiDAR-based 3D object detectors, including PointPillars [30], PV-RCNN [32], and Voxel-RCNN [31], to showcase the versatility of VFMM3D. The impacts of different 3D detectors on prediction accuracy are illustrated in Tab. I.
TABLE II: Ablation analysis on KITTI val set. We quantify the impacts of painting strategy, RGB P means painting all pseudo-LiDAR with corresponding RGB pixel, SAM P refers to pseudo-LiDAR painting by text-prompt SAM.
@@ V Conclusion
In this paper, we have introduced VFMM3D, a novel framework for monocular 3D object detection that leverages Vision Foundation Models (VFMs) to unlock the potential of monocular image data for 3D object detection. By integrating the Segment Anything Model (SAM) and Depth Anything Model (DAM), VFMM3D generates pseudo-LiDAR data that enriches the depth features with detailed spatial information, without the need for being finetuned on specific datasets. Our proposed approach addresses key challenges in monocular 3D object detection, including computational efficiency and noise reduction. VFMM3D could be seamlessly integrated into diverse LiDAR-based detectors, which enhances its versatility and applicability across detection architectures. Through extensive experiments on the KITTI dataset, we have demonstrated the superior performance of VFMM3D over existing methods across various difficulty levels. Our framework consistently outperforms state-of-the-art approaches, showcasing its efficacy in extracting precise 3D object representations from monocular images. The results highlight the potential of VFMM3D as a robust and adaptable solution for real-world deployment in autonomous driving and robotics applications.

### v2

@@ IV-C KITTI Results
We compare the performance of our method on the KITTI val. set for the car class with the state-of-the-art methods. The related results are shown in Tab. II. It is evident that our method outperforms existing methods in both 3D and BEV under all difficulty levels. Specifically, when IoU is set to 0.5, our method surpasses these methods by even greater margins. Compared to the latest monocular object detection method MonoNeRD [53], our approach surpasses it by 8.41% and 12.75% in 3D and BEV at the easy level, respectively. Some qualitative results of 3D detection boxes from our method on KITTI val. set are visualized in Fig. 3.
IV-D Visualization of Each Component Results.
@@ IV-E Ablation Study
Impacts of Different Painting Strategy. We conduct ablation experiments to assess the impacts of different pseudo-LiDAR painting strategies on the final 3D detection results with the PV-RCNN 3D detector. As depicted in Tab. III, we observe that the detailed and accurate foreground information provided by the text-prompt SAM (SAM-P) enables subsequent LiDAR-based 3D detectors to concentrate on the features of foreground objects efficiently. In comparison to the global painting approach (RGB-P), this enhances the overall average detection accuracy by 1.51% and 3.39% for 3D and BEV detection results, respectively.
Ablation Study on Different LiDAR-Based 3D Object Detectors. We employ various LiDAR-based 3D object detectors, including PointPillars [42], PV-RCNN [44], and Voxel-RCNN [43], to showcase the versatility of VFMM3D. The impacts of different 3D detectors on prediction accuracy are illustrated in Tab. II and Tab. I.
TABLE III: Ablation analysis on KITTI val. set. We quantify the impacts of painting strategy, RGB P means painting all pseudo-LiDAR with corresponding RGB pixel, SAM P refers to pseudo-LiDAR painting by text-prompt SAM.
@@ V Conclusion
In this paper, we have introduced VFMM3D, a novel framework for monocular 3D object detection that leverages Vision Foundation Models (VFMs) to unlock the potential of monocular image data for 3D object detection. By integrating the Segment Anything Model (SAM) and Depth Anything Model (DAM), VFMM3D generates pseudo-LiDAR data that enriches the depth features with detailed spatial information without the need for being finetuned on specific datasets. Our proposed approach addresses key challenges in monocular 3D object detection, including computational efficiency and noise reduction. VFMM3D could be seamlessly integrated into diverse LiDAR-based detectors, which enhances its versatility and applicability across detection architectures. Through extensive experiments on the Waymo and KITTI datasets, we have demonstrated the superior performance of VFMM3D over existing methods across various difficulty levels. Our framework consistently outperforms state-of-the-art approaches, showcasing its efficacy in extracting precise 3D object representations from monocular images. The results highlight the potential of VFMM3D as a robust and adaptable solution for real-world deployment in autonomous driving and robotics applications.